OpenStudy (anonymous):

Help please!!!!!!!! someone.... what is the value of x in the kite if TU=15-23 and UV = 3x + 37

OpenStudy (anonymous):

 TU =15-23 and UV =3x +37

OpenStudy (littlebird):

Are you absolutely certain TU=15-23?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

with this figure

OpenStudy (littlebird):

TU is negative. A side length can not be negative.

OpenStudy (anonymous):

yes I know but it is it is maybe because they ask for the value of x

OpenStudy (littlebird):

If you can figure out what TU is you can easily solve the rest of the problem. This diagram should give you a hint.

OpenStudy (anonymous):

oh thanks a lot

OpenStudy (littlebird):

Here's another picture to clarify.

OpenStudy (anonymous):

thanks a lot.